# Leveraging Discrete Function Decomposability for Scientific Design

### v1

@@ A.4 Alternate DADO Derivation: Maximum Entropy Objective
We begin from the maximum entropy problem [jaynes57_mep, ziebart08_maxent, zhu24_aav_max-entropy],
@@
This objective is arguably more consistent with the desired end result for many scientific design problems. Often, a distribution over good solutions is preferred to a single solution (i.e., a distribution with no entropy) [^1] In our experiments (no entropy bonus), the search distribution avoids collapse because we don’t run until convergence (fixed number of iterations), and because we sweep the learning rate (if it’s very high, the distribution may quickly collapse to a point, limiting further improvement). Similar strategies were used by brookes19_cbas instead of an explicit entropy bonus. EDAs were originally used for discrete optimization problems where a single solution was desired. Using distributional optimization to solve more nuanced problems calls for some modifications.. When optimizing a predictive model or more broadly, dealing with uncertainty or inaccuracy in the objective function, having a distribution over good solutions can be particularly important. The maximum entropy objective has a closed-form solution, $p^{*}(x)\propto\exp f(x)/\beta$, motivating the solution of an equivalent variational objective, $\arg\min_{\theta}D_{\mathrm{KL}}\!\left(p_{\theta}\,\|\,p^{*}\right)$, from which we will obtain a decomposed update rule. We first plug in decomposed versions of $f$ and $p_{\theta}(x)$, as described in Sec.Sec. 2.1A.3, to get
@@
To derive the optimal factors, $p^{*}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})$, we’ll use message-passing, for which we now introduce some notation. Note that we adopt the same convention as above of including the root with the rest of the value functions, meaning writing it with a parent, although it doesn’t have a parent and all terms involving its nonexistent parent are simply 0. We recursively define value functions (consistent with levine18_control-as-inference) as:
@@
Because $p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})$ only appears in the KL divergence, the overall expectation will be maximized, as far as $p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})$ is concerned, when the divergence is 0. Therefore, for a leaf node, $p^{*}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})=\exp((Q_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})-V_{i}(\tilde{x}_{\text{p}(i)}))/\beta)$, which is a proper distribution because $\exp V_{i}(\tilde{x}_{\text{p}(i)})$ is exactly the normalizing constant of $\exp Q_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})$. Notice that we still have a term, $\mathbb{E}_{p_{\theta}(\tilde{x}_{\text{p}(i)})}\left[{V_{i}(\tilde{x}_{\text{p}(i)})}/{\beta}\right]$, outside of the divergence (Eq. 48). Maximizing this term must therefore be the responsibility of $p_{\theta}(\tilde{x}_{\text{p}(i)})$, not $p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})$. In fact, these leftover terms are included in the parent’s $Q$-function via the sum over all child $V$-functions in $Q_{i}(\tilde{x}_{i})$ as we defined it above (Eq. 46). It follows that the recursive case has the same form as the leaves, which had $Q_{i}(\tilde{x}_{i})=f_{i}(\tilde{x}_{i})$ because they’ve no children (more detailed derivation in levine18_control-as-inference). That is, for all non-root nodes $i$, $p^{*}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})=\exp((Q_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})-V_{i}(\tilde{x}_{\text{p}(i)}))/\beta)$; for the root node, $p^{*}(\tilde{x}_{r})=\exp((Q_{r}(\tilde{x}_{r})-V_{r})/\beta)$, where $V_{r}=\log\sum_{\tilde{x}_{r}}\exp Q_{r}(\tilde{x}_{r})$ is a normalizing constant with no dependence on $\theta$. We now have an optimal form for each factor and can proceed to minimize our search factors’ divergences from them completely separately.
@@
There are a few minor differences from DADO. First, there’s an entropy-maximizing term. Second, notice that the value function definitions are a little different from in DADO. Here we use $\log\sum\exp$ to define $V_{i}(\tilde{x}_{\text{p}(i)})$, whereas DADO uses an expectation over the search distribution instead. Third, the $Q$-function has $V_{i}(\tilde{x}_{\text{p}(i)})$ subtracted from it, whereas DADO’s derivation doesn’t. As mentioned above though, $V_{i}(\tilde{x}_{\text{p}(i)})$’s role in the objective is that of a “baseline”, and it can be changed out or dropped. In our implementation, we do subtract $V_{i}(\tilde{x}_{\text{p}(i)})$—the expectation-based one, not $\log\sum\exp$—as a mean baseline function to reduce weight variance and obtain a more stable update (Sec. 4). Overall though, the result is quite similar to DADO (Sec. A.3), supporting the generality of the decomposition-aware distributional optimization framework, and suggesting that there are a variety of related objectives and algorithms one might use under this umbrella. In particular, one might adapt methods from RL. For example, peng19_awr’s derivation could be adapted to get a similar weighted maximum likelihood objective with the difference in value functions (also called the advantage) exponentiated instead. TRPO, PPO, or CbAS [schulman15_trpo, schulman17_ppo, brookes19_cbas] might be adapted to regularize the search distribution toward some (identically factorized) prior, or the previous iteration’s search distribution for stability. We emphasize that the core idea of DADO is to leverage message-passing (i.e., value functions) on arbitrarily decomposed objectives to derive a distributional optimization procedure which updates search distribution factors separately, regardless of the specific objective and algorithm used.
@@
A.5 Protein Dataset Details
@@
We chose four protein property datasets based on three being readily available in a convenient form from Mave2022, and one from notin23_proteingym that has sequences many mutations away from the wild-type assayed. These proteins span various lengths and all have more than single mutations away from wild-type. These datasets are:
  • Amyloid-beta (Amyloid; $L=42,D=21$, where the extra state is a stop codon), with data from seuma21_amyloid_beta, which assayed 16,066 sequences for aggregation with a nucleation score, 97% of which are double mutants. The dataset was accessed via Mave2022.
@@
  • Adeno-associated virus 2 capsid protein (AAV2 capsid or AAV; $L=28,D=20$), with data from bryant21_aav, which assayed 42,329 sequences for virus viability, including sequences with as many as 27 mutations from the wild-type. The full protein is of length 735. The dataset was accessed via notin23_proteingym.
  • Protein G B1 domain (GB1; $L=55,D=20$), with data from olson14_gb1, who scanned all possible single and double mutants and assayed their 536,963 binding affinities to IgGFC, of which 99.8% are double mutants. The dataset was accessed via Mave2022.
@@
  • TAR DNA-binding protein 43 (TDP-43; $L=84,D=21$, where the extra state is a stop codon), with data  bolognesi19_tdp43, which assayed 57,996 sequences for cell toxicity, 98% of which are double mutants. The dataset was accessed via Mave2022.
We considered a few other datasets only shown in experiments in the appendix (Sec. A.6.2). These were also chosen from  notin23_proteingym on the basis of containing sequences many mutations away from the wild-type assayed, plus one arbitrarily chosen from the shortest-length (for expediency) domains in  tsuboyama23_mega. These datasets are:
  • ynzC, a small protein domain ($L=39,D=20$), with data from tsuboyama23_mega, assayed for its folding stability on 2,301 variants, 68% of which are double mutants. The dataset was accessed via notin23_proteingym.
  • Chlamydomonas reinhardtii’s light-oxygen-voltage domain (CreiLOV or Phot; $L=118,D=20$), with data from chen23_phot, which assayed 167,530 sequences for fluorescence, including sequences with as many as 15 mutations from the wild-type. The dataset was accessed via notin23_proteingym.
@@
  • Yeast transcription factor Gcn4 ($L=44,D=20$), with data from staller18_gcn4, which assayed 2,639 sequences for activity, including sequences with as many as 44 mutations from the wild-type. The full protein is of length 281. The dataset was accessed via notin23_proteingym.
It’s worth noting that the Gcn4 activation domain assayed is intrinsically disordered, but folds into a helix when it binds its coactivator [staller18_gcn4]. We expect that our rudimentary procedure of obtaining a contact map from an AlphaFold3 [abramson24_af3]) 3D structure won’t work for proteins that exist in more than one stable configuration because the resulting decomposition will not capture interactions crucial to the assayed property. We find that the resulting predictive model has very poor holdout accuracy. Developing more accurate decompositions for such problems is an interesting line of future work.
A.6 Additional experimental results
Here, we include additional experimental results for completeness. For synthetic landscapes and for protein landscapes, we ablate the number of samples drawn at each EDA iteration compared to the main text figures. For protein landscapes, we also perform a case study on one protein dataset in order to study the effect of the distance threshold, $t$, on decomposability, predictive accuracy, and downstream optimization performance.
A.6.1 Synthetic functions
We first present an ablation on number of samples drawn each iteration, $K$, between DADO and a decomposition-unaware EDA on synthetic functions, keeping everything else the same as in Sec. A.6.1. For this ablation, we only consider synthetic landscapes with sequence lengths of $L=25$, $L=50$, $L=100$; we omit $L=200$ (Fig. 2c) for expediency. Optimization results for both $K=100$ (Fig. A1) and $K=1000$ (Fig. A2) support the conclusion that DADO is able to find designs with higher $f(x)$ earlier than the EDA, and by the end of 100 iterations, has a comparable or larger sample mean than the EDA. Both methods, as one would expect when performing distributional optimization, achieve faster and more effective optimization when the sampling budget is increased. DADO shows small performance improvements when given tenfold samples, suggesting that it is pretty sample-efficient and has already reached saturating performance by $K=100$, whereas the EDA finds samples with substantially higher $f(x)$ when given the same increased sample budget, demonstrating that it had not yet saturated at $K=100$.
[…]
Figure A1: Comparison of a standard EDA to DADO on synthetic problems drawing 100 samples each iteration. We created three random functions, $f(x)$, each with a randomly chosen junction tree decomposition with maximum node size of one, and randomly chosen parameters. Each experiment used alphabet size $D=20$ and sequence length, a, $L=25$ (also Fig. 2a), b, $L=50$ (also Fig. 2b), and c, $L=100$. Each of the two methods drew $K=100$ samples per iteration. For each iteration, we show the mean (solid line) and 95% confidence interval (shaded envelope) of the 100 samples evaluated on $f(x)$, averaged across results from 20 random seeds. P-values are from two-sided paired t-tests that AUC of the per-iteration mean is different between methods, using the 20 mean curves.
[…]
Figure A2: Comparison of a standard EDA to DADO on synthetic problems drawing 1000 samples each iteration. We created three random functions, $f(x)$, each with a randomly chosen junction tree decomposition with maximum node size of one, and randomly chosen parameters. Each experiment used alphabet size $D=20$ and sequence length, a, $L=25$, b, $L=50$, and c, $L=100$. Each of the two methods drew $K=1000$ samples per iteration. For each iteration, we show the mean (solid line) and 95% confidence interval (shaded envelope) of the 100 samples evaluated on $f(x)$, averaged across results from 20 random seeds. P-values are from two-sided paired t-tests that AUC of the per-iteration mean is different between methods, using the 20 mean curves.
A.6.2 Protein property predictors
We now present an ablation on seven protein landscapes for the number of samples at each iteration, $K$, keeping all other settings the same as in Sec. 4.2. These include the four proteins originally shown in the main text—Amyloid, AAV, GB1, and TDP-43 (Fig. 3)—as well as three proteins not shown elsewhere—ynzC, CreiLOV, and Gcn4—so that all can be viewed simultaneously and compared. The datasets upon which each of these protein property predictors is based are described in Sec. A.5.
We first present experimental results on the seven synthetic landscapes with $K=1000$ (Fig. A3). For this setting, optimization performance for both DADO and the EDA is saturated and comparable for two of the additional proteins, ynzC and CreiLOV (Fig. A3e,f). On Gcn4, for which the decomposed predictive model is based on the structure of an intrinsically disordered protein and so does not fit the assay very well (details in Sec. A.5), DADO is able to find high $f(x)$ designs more effectively than the EDA. All other predictive models acheived good holdout accuracy; we include Gcn4 primarily as an example of a protein lacking a good decomposition.

### v2

@@ A.5 Alternative DADO Derivation: Maximum Entropy Objective
We begin from the maximum entropy problem (Jaynes, 1957; Ziebart et al., 2008; Zhu et al., 2024),
@@
This objective is arguably more consistent with the desired end result for many scientific design problems. Often, a distribution over good solutions is preferred to a single solution (i.e., a distribution with no entropy) [^1] In our experiments (no entropy bonus), the search distribution avoids collapse because we don’t run until convergence (fixed number of iterations), and because we sweep the learning rate (if it’s very high, the distribution may quickly collapse to a point, limiting further improvement). Similar strategies were used by Brookes et al. (2019) instead of an explicit entropy bonus. EDAs were originally used for discrete optimization problems where a single solution was desired. Using distributional optimization to solve more nuanced problems calls for some modifications.. When optimizing a predictive model or more broadly, dealing with uncertainty or inaccuracy in the objective function, having a distribution over good solutions can be particularly important. The maximum entropy objective has a closed-form solution, $p^{*}(x)\propto\exp f(x)/\beta$, motivating the solution of an equivalent variational objective, $\arg\min_{\theta}D_{\mathrm{KL}}\!\left(p_{\theta}\,\|\,p^{*}\right)$, from which we will obtain a decomposed update rule. We first plug in decomposed versions of $f$ and $p_{\theta}(x)$, as described in Sec. 2.1/Sec. A.4, to get
@@
To derive the optimal factors, $p^{*}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})$, we’ll use message-passing, for which we now introduce some notation. Note that we adopt the same convention as above of including the root with the rest of the value functions, meaning writing it with a parent, although it doesn’t have a parent and all terms involving its nonexistent parent are simply 0. We recursively define value functions (consistent with Levine (2018)) as:
@@
Because $p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})$ only appears in the KL divergence, the overall expectation will be maximized, as far as $p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})$ is concerned, when the divergence is 0. Therefore, for a leaf node, $p^{*}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})=\exp((Q_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})-V_{i}(\tilde{x}_{\text{p}(i)}))/\beta)$, which is a proper distribution because $\exp V_{i}(\tilde{x}_{\text{p}(i)})$ is exactly the normalizing constant of $\exp Q_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})$. Notice that we still have a term, $\mathbb{E}_{p_{\theta}(\tilde{x}_{\text{p}(i)})}\left[{V_{i}(\tilde{x}_{\text{p}(i)})}/{\beta}\right]$, outside of the divergence (Eq. 39). Maximizing this term must therefore be the responsibility of $p_{\theta}(\tilde{x}_{\text{p}(i)})$, not $p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})$. In fact, these leftover terms are included in the parent’s $Q$-function via the sum over all child $V$-functions in $Q_{i}(\tilde{x}_{i})$ as we defined it above (Eq. 37). It follows that the recursive case has the same form as the leaves, which had $Q_{i}(\tilde{x}_{i})=f_{i}(\tilde{x}_{i})$ because they’ve no children (more detailed derivation in Levine (2018)). That is, for all non-root nodes $i$, $p^{*}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})=\exp((Q_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})-V_{i}(\tilde{x}_{\text{p}(i)}))/\beta)$; for the root node, $p^{*}(\tilde{x}_{r})=\exp((Q_{r}(\tilde{x}_{r})-V_{r})/\beta)$, where $V_{r}=\log\sum_{\tilde{x}_{r}}\exp Q_{r}(\tilde{x}_{r})$ is a normalizing constant with no dependence on $\theta$. We now have an optimal form for each factor and can proceed to minimize our search factors’ divergences from them completely separately.
@@
There are a few minor differences from DADO. First, there’s an entropy-maximizing term. Second, notice that the value function definitions are a little different from in DADO. Here we use $\log\sum\exp$ to define $V_{i}(\tilde{x}_{\text{p}(i)})$, whereas DADO uses an expectation over the search distribution instead. Third, the $Q$-function has $V_{i}(\tilde{x}_{\text{p}(i)})$ subtracted from it, whereas DADO’s derivation doesn’t. As mentioned above though, $V_{i}(\tilde{x}_{\text{p}(i)})$’s role in the objective is that of a “baseline”, and it can be changed out or dropped. In our implementation, we do subtract $V_{i}(\tilde{x}_{\text{p}(i)})$—the expectation-based one, not $\log\sum\exp$—as a mean baseline function to reduce weight variance and obtain a more stable update (Sec. 4). Overall though, the result is quite similar to DADO (Sec. A.4), supporting the generality of the decomposition-aware distributional optimization framework, and suggesting that there are a variety of related objectives and algorithms one might use under this umbrella. In particular, one might adapt methods from RL. For example, Peng et al. (2019)’s derivation could be adapted to get a similar weighted maximum likelihood objective with the difference in value functions (also called the advantage) exponentiated instead. TRPO, PPO, or CbAS (Schulman et al., 2015; 2017; Brookes et al., 2019) might be adapted to regularize the search distribution toward some (identically factorized) prior, or the previous iteration’s search distribution for stability. We emphasize that the core idea of DADO is to leverage message-passing (i.e., value functions) on arbitrarily decomposed objectives to derive a distributional optimization procedure which updates search distribution factors separately, regardless of the specific objective and algorithm used.
@@
A.6 Protein Dataset Details
@@
We investigated seven protein property datasets from  Tareen et al. (2022) and Notin et al. (2023). These proteins span various lengths and all have more than single mutations away from wild-type. These datasets are:
@@
  • Adeno-associated virus 2 capsid protein (AAV2 capsid or AAV; $L=28,D=20$), with data from Bryant et al. (2021), which assayed 42,329 sequences for virus viability, including sequences with as many as 27 mutations from the wild-type. The full protein is of length 735. The dataset was accessed via Notin et al. (2023).
@@
  • Amyloid-beta (Amyloid; $L=42,D=21$, where the extra state is a stop codon), with data from Seuma et al. (2021), which assayed 16,066 sequences for aggregation with a nucleation score, 97% of which are double mutants. The dataset was accessed via Tareen et al. (2022).
@@
  • Yeast transcription factor Gcn4 ($L=44,D=20$), with data from Staller et al. (2018), which assayed 2,639 sequences for activity, including sequences with as many as 44 mutations from the wild-type. The full protein is of length 281. The dataset was accessed via Notin et al. (2023).
  • TAR DNA-binding protein 43 (TDP-43; $L=84,D=21$, where the extra state is a stop codon), with data  Bolognesi et al. (2019), which assayed 57,996 sequences for cell toxicity, 98% of which are double mutants. The dataset was accessed via Tareen et al. (2022).
  • Protein G B1 domain (GB1; $L=55,D=20$), with data from Olson et al. (2014), who scanned all possible single and double mutants and assayed their 536,963 binding affinities to IgGFC, of which 99.8% are double mutants. The dataset was accessed via Tareen et al. (2022).
  • ynzC, a small protein domain ($L=39,D=20$), with data from Tsuboyama et al. (2023), assayed for its folding stability on 2,301 variants, 68% of which are double mutants. The dataset was accessed via Notin et al. (2023).
  • Chlamydomonas reinhardtii’s light-oxygen-voltage domain (CreiLOV or Phot; $L=118,D=20$), with data from Chen et al. (2023), which assayed 167,530 sequences for fluorescence, including sequences with as many as 15 mutations from the wild-type. The dataset was accessed via Notin et al. (2023).
Of these, the four shown in the main text were chosen according to two criteria. First, we wanted proteins for which the AlphaFold3-derived junction tree had relatively small nodes (i.e., more decomposable proteins), discussed in more detail below. Second, we also prioritized proteins for which the datasets contained sequences many mutations away from the wild-type, as the resulting predictive models are more likely to be realistic in a larger area of the design space than datasets that only assay a concentrated ball of sequences. Based on these criteria, we decided to focus on AAV, Amyloid, Gcn4, and TDP-43 (Fig. 3). These four proteins’ decomposition junction trees have nodes of cardinality five or fewer, whereas the junction trees for GB1 and ynzC have many nodes with cardinality over 10, and CreiLOV has nodes with cardinality greater than 20. Optimization performance on the remaining datasets is included in the appendix (Fig.A7).
We visualize all seven proteins’ AlphaFold3-predicted structures and corresponding contact graphs and junction trees (Fig. A4). Panels 3–6 (the rightmost four columns) illustrate that AAV, Amyloid, Gcn4, and TDP-43 yield junction trees with smaller nodes and fewer positions overlapping in multiple junction tree nodes compared to GB1, ynzC, and CreiLOV. It would be interesting to reduce the contact threshold for the latter problems in order to obtain smaller nodes. As the possible efficiency gain from using a decomposition-aware method like DADO is determined by how sparse and tree-like this graph is, we expect DADO would yield more substantial performance improvements over standard EDAs in such a setting. Our studies of predictive model accuracy when varying the contact threshold (Figs. A9A10A12) suggest that one may be able to obtain almost-as-accurate predictive models under decomposition graphs with lower contact thresholds, but we did not further investigate optimization performance on these models.
@@
[…]
Figure A4: 3D structure, contact graph, and junction tree decomposition for each protein. For each of the seven protein datasets considered, a, AAV; b, Amyloid-$\beta$; c, Gcn4; d, TDP-43; e, GB1; f, ynzC; and g, CreiLOV, we show the AlphaFold3-predicted 3D structure and the contact graph and junction tree (JT) decomposition obtained by imposing a $4.5\,\AA$ contact threshold. From left to right, each panel shows: (1) the full 3D backbone structure, with the $\alpha$-carbon atoms of designable positions highlighted in red; (2) the contact graph restricted to designable positions, where edges connect residue pairs with any atoms within $4.5\,\AA$ of each other; (3) positions in the contact graph colored by how many JT nodes each participates in (global scale 1–25); (4) a 2D PCA projection of positions overlaid with convex hulls representing each JT node (hulls may encompass positions not contained as a result of dimensionality reduction), colored by cardinality (global scale 1–26) and with the root node marked by a thick black outline; (5) the directed JT topology, with nodes colored by cardinality (global scale 1–26) and the root node marked by a thick black outline; and (6) each node’s cardinalities (top) and edge cardinalities—that is, of the intersection of the node’s variables with its parent’s variables—(bottom) with shared global scale (1–52). To examine high-resolution details (e.g., 3D residue orientations), please view on a screen and zoom in.
A.7 Baseline method details
We compare DADO with three baseline methods. The first is the naive, decomposition-unaware EDA, which we refer to as “EDA”. Algorithmic details of this method are in Sec. 2 and Sec. A.2 also includes a derivation of the EDA algorithm. We take this as the starting point for the other baseline methods and now describe how the EDA is modified for each.
The second baseline we consider is the factorized distribution algorithm (“FDA”) from Mühlenbein and Mahnig (1999). The only modification to the naive EDA is to replace its search distribution, ordinarily a joint distribution over all design variables, with a factorized search distribution. In all of our experiments, we give FDA the same search distribution factorization as DADO for fairness; as such, in our protein experiments, FDA’s factorization is based on the AlphaFold3-based contact map. Comparing to FDA sheds light on how much of DADO’s improved performance can be attributed to it using a factorized search distribution versus the contribution of using message-passing value functions for the factorized search distribution update because FDA only uses the former and not the latter, whereas DADO uses both.
The third baseline we consider is a proximal policy optimization version of the EDA, which we call “PPO”. This baseline is inspired by Schulman et al. (2017). We only modify the EDA’s update rule to match the PPO update rule. That is, instead of minimizing the naive EDA’s loss, $L(\theta)=\mathbb{E}_{x^{k}\sim p_{\theta^{n}}(x)}\left[f(x^{k})\log p_{\theta}(x^{k})\right]$, we instead minimize a proximal version of it,
$L(\theta)=\mathbb{E}_{x^{k}\sim p_{\theta^{n}}(x)}\left[\text{min}\left\{f(x^{k})\frac{p_{\theta}(x^{k})}{p_{\theta^{n}}(x)},f(x^{k})\;\text{clip}(\frac{p_{\theta}(x^{k})}{p_{\theta^{n}}(x)},1-\epsilon,1+\epsilon)\right\}\right],$
where $\epsilon$ is a hyperparameter that we set to $0.2$ as suggested in the original PPO paper. PPO-style updates can be interpreted as stabilizing the search distribution update of the naive EDA by penalizing it from deviating too far from the previous iteration’s search distribution. This approach could easily be added onto DADO and FDA; we did not explore this but expect it would be helpful for these decomposition-aware methods too.
A.8 Additional Implementation Details
We record various implementation details for reproducibility (also see code released upon publication). We implemented all of our code using Jax, Flax, and Optax (DeepMind et al., 2020).
A.8.1 Search distribution
For the search distribution of DADO, we used MLP-based autoregressive neural networks for each position $l\in[1,\ldots,L]$—specifically, an MLP of size $[64,64]$ that takes as input all conditioning variables (parents in the tree) and outputs $D$ logits for each step of autoregressive decoding. Our naive EDA implementation is the same, only assuming all variables are in one meta-variable—hence, a standard autoregressive model over all variables. The FDA baseline uses the same architecture as DADO, whereas our PPO baseline uses the same architecture as the naive EDA. For gradient descent to fit the search distribution, we used the AdamW optimizer (Loshchilov and Hutter, 2017) with default momentum parameters $\beta_{1}=0.9,\beta_{2}=0.999$.
A.8.2 Fully synthetic functions
We simulated one $f(x)$ for each of three sequence lengths, $L=\{25,50,200\}$, by first randomly sampling a junction tree topology, and then randomly specifying the component functions. Each node function, a $D$-vector, was sampled $f_{i}\in\mathbb{R}^{D}\sim\mathcal{N}(0,0.01)$, where $\mathcal{N}$ denotes a Gaussian distribution. Each edge function, a $D\times D$ matrix, was sampled, $f_{i,j}\in\mathbb{R}^{D\times D}\sim\mathcal{N}(0,0.0025)$. To ensure a reasonable degree of non-smoothness in $f(x)$, we further explicitly added what in biology is known as reciprocal sign epistasis (Starr and Thornton, 2016; Li et al., 2024). Specifically, for each edge function we, twice, randomly assigned one of the 20 alphabet letters to each node, i) $x_{i}:=A,x_{j}:=B$ and ii) $x_{i}:=C,x_{j}:=D$. Next we sampled an effect size, $\lambda\sim\mathcal{N}(0,4)$. Finally, we set $f_{i,j}(x_{i}=A,x_{j}=B)=0$ and $f_{i,j}(x_{i}=C,x_{j}=B)=\lambda$, and also half of the time, $f_{i,j}(x_{i}=A,x_{j}=D)=0$ and $f_{i,j}(x_{i}=A,x_{j}=D)=\lambda$.
A.8.3 Decomposed protein property predictive models
For each dataset, we used AlphaFold3-predicted structures (Abramson et al., 2024) on the wild-type sequence to obtain a 3D structure, from which we constructed a contact graph by thresholding the distance between pairs of residues with threshold $t$. Following Brookes et al. (2022); Romero et al. (2013); Voigt et al. (2002), we use a threshold of $t=4.5\text{\AA }$. We interpret this contact map as a graph adjacency matrix, from which we algorithmically construct a junction tree (Lauritzen and Spiegelhalter, 1988). This defines the topology needed for DADO, but we must also fit the component functions on the protein assay-labeled data. To do so, we employ an MLP-based predictive model that strictly enforces the decomposition defined by the junction tree. In particular, we use an MLP that takes as input the sequence, and, critically, also a bit-vector specifiying active variables. One function evaluation requires calling this MLP for every node and edge function and summing. We used 5-fold cross validation to sweep through hidden layers of size ($[16,16]$, $[128,16]$, or $[128,128,16]$), learning rate ($0.001$ or $0.0001$), and number of training iterations (5,000 or 50,000), to choose hyperparameters with the lowest cross-validation mean-squared error. Finally, we trained the model with those hyper-parameters using all data available.
A.9 Additional Experimental Results
Herein, we include additional experimental results on synthetic landscapes for completeness—primarily studying the relationship between performance and number of design variables ($L$), and larger alphabet sizes ($D$). For protein landscapes, we also perform a case study on one protein dataset in order to study the effect of the distance threshold, $t$, on decomposability, predictive accuracy, and downstream optimization performance.
A.9.1 Synthetic functions: L=400
We wanted to compare DADO and a naive EDA on even larger design spaces, but this can take quite a while to run when considering the hyperparameter sweep and replicates (Sec. A.9.7). For expediency, we chose the same hyperparameters (temperature and learning rate) as were used by each method for $L=200$ (Fig. 2). We then ran 20 replicates of each method.
At $L=400$, the trend we observed in Fig. 2 continues—as $L$ grows, the gap between DADO and the EDA grows increasingly larger (Fig. A5), as expected, because the full design space grows at a faster rate than the decomposed design space that DADO operates in.